1. The cost, in millions of dollars, to remove x % of pollution in a lake modeled by C = 6,000/200-2(X)
a. What is the cost to remove 75% of the pollutant?
b. What is the cost to remove 90 % of the pollutant?
c. What is the cost to remove 99% of the pollutant?
d. For what value is this equation undefined?

Answer by checkley77(12844)

75%,90%,95%,undefined?
C = 6,000/200-2(X)
C=6,000/200-2*75
C=6,000/200-150
C=6,000/50
C=120 MILLION.
C=6,000/200-2*90
C=6,000/200-180
C=6,000/20
C=300 MILLION.
C=6,000/200-2*95
C=6,000/200-190
C=6,000/10
C=600 MILLION.
100% RESULTS IN AN UNDEFINED COST.
C=6,000/200-2*100
C=6,000/200-200
C=6,000/0
DIVIDING BY 0 RESULTS IN AN UNDEFINED ANSWER.